A lot of 12 television sets chosen at random are defective , if 3 of the sets are chosen at random for shipment in hotel. how many defective set can they expect
1
Expert's answer
2020-06-11T16:44:03-0400

Let x denote the number of defective televisions out of the 12. The probability(p) of selecting a defective television is "\frac {x} {12}"

Mean =np;

="3*\frac {x} {12}=\frac {x} {4}"

They should expect "\frac {x} {4}" defective sets